Question Help * lf a lender makes a simple loan of $70 or 3 years and charges 7%, then the amount that the lender receive at maturity s S Round your esponse o the nearest two decimal place) If a lender makes a simple loan of $3000 for one year and charges $70 interest, then the simple interest rate on that loan is%. (Round your response to the nearest whole number) If a borrower must repay $106.50 one year from today in order to receive a simple loan of $100 today, the simple interest on this loan is OA. 65% B. 6.0% O c. 5.090 D. 6.590 Click to select your answer(s). oly Relic From Two Thousa Years Ago 2002 No thanks Try it free
Answer #1

Answer:

A) Loan amount P=$700

Interest rate r=7%

Number of years = 3years

So amount lender receives at maturity= 3*7%*700+700=$847

B)

Loan amount P=$3000

Interest I=70

Number of years = 1 years

Simple interest rate r=70/(3000*1)=2.33%

C)

Loan amount P=$100

Borrower pays at end of 1 year A=$106.5

Interest I=A-P=106.5-100=6.5

Number of years = 1 years

Simple interest rate r=6.5/(100*1)=6.5%